OpenStudy (anonymous):

Derivative 1-cos^2(2x)

OpenStudy (anonymous):

use the chain rule: d/dx(f(g(x))=f'(g(x))*g'(x) f(x)=x^2 g(x)=cos(2x) d/dx (cos^2(2x)) = 2cos(2x)*(-2 sin(2x))